# Impact of existence and nonexistence of pivot on the coverage of empirical best linear prediction intervals for small areas

### v2

@@ 3 Single parametric bootstrap
$\begin{split}&\text{E}\left[(\theta_{i}-\tilde{\theta}_{i})/\sqrt{g_{1i}}\right]^{4}\\ &=\frac{1}{g_{1i}^{2}}\text{E}\left[\bm{x_{i}^{\prime}\beta}+u_{i}-(1-B_{i})y_{i}-B_{i}\bm{x_{i}^{\prime}\beta}\right]^{4}\\ &=\frac{1}{g_{1i}^{2}}\text{E}\left[B_{i}u_{i}-(1-B_{i})e_{i}\right]^{4}\\ &=\frac{1}{g_{1i}^{2}}\text{E}\left[B_{i}^{4}u_{i}^{4}-4B_{i}^{3}u_{i}^{3}(1-B_{i})e_{i}+6B_{i}^{2}u_{i}^{2}(1-B_{i})^{2}e_{i}^{2}-4B_{i}u_{i}(1-B_{i})^{3}e_{i}^{3}+(1-B_{i})^{4}e_{i}^{4}\right]\\ &=\frac{1}{g_{1i}^{2}}\text{E}\left[B_{i}^{4}u_{i}^{4}+6B_{i}^{2}u_{i}^{2}(1-B_{i})^{2}e_{i}^{2}+(1-B_{i})^{4}e_{i}^{4}\right]\\ &=3+\frac{D_{i}^{2}}{(A+D_{i})^{2}}\left[\frac{\text{E}(u_{i}^{4})}{A^{2}}-3\right],\end{split}$ (3.6)
@@ 4 Double parametric bootstrap
Hall and Maiti (2006) considered parametric bootstrap methods to approximate the distribution of $(\theta_{i}-\bm{x_{i}^{\prime}\hat{\beta}})/\sqrt{\hat{A}}$. Then, the prediction interval can be constructed as $(\bm{x_{i}^{\prime}}\hat{\bm{\beta}}+b_{\hat{\alpha}_{l}^{s}}\sqrt{\hat{A}},\bm{x_{i}^{\prime}}\hat{\bm{\beta}}+b_{\hat{\alpha}_{u}^{s}}\sqrt{\hat{A}})$, where $b_{\hat{\alpha}_{l}^{s}}$ and $b_{\hat{\alpha}_{u}^{s}}$ are obtained from single bootstrap approximation based on their algorithm. Their prediction interval is based on a synthetic model or the regression model, which does not permit approximation of the conditional distribution of $\theta_{i}$ given the data $\bm{Y}$. As a consequence, it is likely to underweight the area specific data. When the level 2 distribution $\mathcal{G}$ determined only by $\bm{x_{i}^{\prime}\beta}$ and $A$, it is easy to know that the quantity, $(\theta_{i}-\bm{x_{i}^{\prime}\beta})/\sqrt{A}$, is a pivot. As Hall and Maiti (2006) stated, their prediction interval is effective as $\mathrm{P}(\bm{x_{i}^{\prime}}\hat{\bm{\beta}}+b_{\hat{\alpha}_{l}^{s}}\sqrt{\hat{A}}\leq\theta_{i}\leq\bm{x_{i}^{\prime}}\hat{\bm{\beta}}+b_{\hat{\alpha}_{u}^{s}}\sqrt{\hat{A}})=1-\alpha+O(m^{-2})$. However, when additional parameters involved into the distribution of random effects, $(\theta_{i}-\bm{x_{i}^{\prime}\beta})/\sqrt{A}$ might not a pivot and we may lose the effectiveness.
Proposition 4.1.
@@
$\mathrm{P}(\bm{x_{i}^{\prime}}\hat{\bm{\beta}}+b_{\hat{\alpha}_{l}^{s}}\sqrt{\hat{A}}\leq\theta_{i}\leq\bm{x_{i}^{\prime}}\hat{\bm{\beta}}+b_{\hat{\alpha}_{u}^{s}}\sqrt{\hat{A}})=1-\alpha+O(m^{-1}),$ (4.1)
@@
See proof in the appendix D, which is similar to Proposition 3.1.
@@
While the order $O(m^{-1})$ term $T_{1}$ in (3.5) is theoretically positive under certain conditions, it remains unclear whether this positiveness holds when $u_{i}$ is asymmetrically distributed. Unlike Hall and Maiti (2006), we introduce a new double bootstrap method, which does not require a pivot or symmetric $u_{i}$. This method reduces the coverage error to $o(m^{-1})$, even when $u_{i}$ is asymmetrically distributed. Our double bootstrap approach is based on the algorithm from Shi (1992), where the double bootstrap is proposed to obtain accurate and efficient confidence interval for parameters of interest in both nonparametric and univariate parametric distribution settings. Later on, McCullough and Vinod (1998) discussed the theory of the double bootstrap both with and without pivoting, and provided the implementations in some nonlinear production functions. In this paper, we develop the double bootstrap in the context of our mixed effect model and apply it to obtain the EBL prediction intervals of $\theta_{i}$. The framework of our double parametric bootstrap is as below:
@@
      Conditionally on the data $(\bm{X,Y})$, draw $\theta_{i}^{*},\,i=1,\cdots,m$, from the distribution $\mathcal{G}(\bm{x_{i}}^{\prime}\hat{\bm{\beta}},\hat{A},\hat{\bm{\phi}})$;
@@
      Compute $\hat{\bm{\beta}}^{*}$, $\hat{A}^{*}$, $\hat{\bm{\phi}}^{*}$ from the data $(\bm{X,Y^{*}})$, and obtain $\hat{\theta}_{i}^{*}$ and $\hat{g}_{1i}^{*}$;
@@
      Compute $\hat{\bm{\beta}}^{**}$, $A^{**}$ from the data $(\bm{X,Y^{**}})$, also obtain $\hat{\theta}_{i}^{**}$ and $\hat{g}_{1i}^{**}$;
@@
      $\hat{G}^{*}(z)\equiv\mathrm{P}^{**}(H_{i}(\hat{\bm{\beta}}^{**},\hat{A}^{**})\leq z\,|\,\hat{\bm{\beta}}^{*},\hat{A}^{*},\hat{\phi}^{*})$ (4.2)
@@
      $\left\{\begin{array}[]{@{}l@{}}\mathrm{P}^{*}(H_{i}(\hat{\bm{\beta}}^{*},\hat{A}^{*})\leq q_{\hat{\alpha}_{u}}\,|\,\hat{\bm{\beta}},\hat{A},\hat{\bm{\phi}})=1-\alpha/2\\ \mathrm{P}^{**}(H_{i}(\hat{\bm{\beta}}^{**},\hat{A}^{**})\leq q_{\hat{\alpha}_{u}}\,|\,\hat{\bm{\beta}}^{*},\hat{A}^{*},\hat{\bm{\phi}}^{*})={\hat{\alpha}_{u}}.\end{array}\right.\,,$ (4.3)
      Using the definition (4.2), we have $q_{\hat{\alpha}_{u}}=\hat{G}^{*-1}(\hat{\alpha}_{u})$. Then, the above system of equations is equivalent to
      $\mathrm{P}^{*}\left(H_{i}(\hat{\bm{\beta}}^{*},\hat{A}^{*})\leq\hat{G}^{*-1}(\hat{\alpha}_{u})\,|\,\hat{\bm{\beta}},\hat{A},\hat{\bm{\phi}}\right)=1-\alpha/2.$ (4.4)
      Rewriting (4.4) gives
      $\mathrm{P}^{*}\left[\mathrm{P}^{**}\left(H_{i}(\hat{\bm{\beta}}^{**},\hat{A}^{**})\leq H_{i}(\hat{\bm{\beta}}^{*},\hat{A}^{*})\,|\,\hat{\bm{\beta}}^{*},\hat{A}^{*},\hat{\bm{\phi}}^{*}\right)\leq\hat{\alpha}_{u}\,|\,\hat{\bm{\beta}},\hat{A},\hat{\bm{\phi}}\right]=1-\alpha/2.$ (4.5)
@@
      We will use $Z_{j}$ to adjust the first-stage intervals. After all bootstrapping operations are complete, we have estimates $Z_{j}$, $j=1,2,\cdots,J$, where $J$ is the number of first-stage bootstrap operations. Sort $Z_{j}$ and choose the $21-\alpha/$ 2 quantile as the percentile point $\hat{\alpha}_{u}$ for defining the double-bootstrap upper limit.
@@
      $\mathrm{P}^{*}(H_{i}(\hat{\bm{\beta}}^{*},\hat{A}^{*})\leq q_{\alpha_{u}}\,|\,\hat{\bm{\beta}},\hat{A},\hat{\bm{\phi}})=\hat{\alpha}_{u}.$ (4.6)
@@
The algorithm above shows that the single parametric bootstrap $\hat{I}^{s}_{i}$ is calibrated by the second-stage bootstrap. In general, such a calibration improves the coverage accuracy to $o(m^{-1})$, even when the pivot does not exist. One more advantage of our double bootstrap algorithm is that it avoids the problem of over correction and so make it more practical than that of Hall and Maiti (2006).
Theorem 4.1.
@@
$\mathrm{P}\left(\hat{\theta}_{i}+q_{\hat{\alpha}^{d}_{l}}\sqrt{\hat{g}_{1i}}\leq\theta_{i}\leq\hat{\theta}_{i}+q_{\hat{\alpha}_{u}^{d}}\sqrt{\hat{g}_{1i}}\right)=1-\alpha+o(m^{-1}),$ (4.7)
@@
See proof of the theorem in appendix E.
@@ 5 Monte Carlo Simulations
In this section, we compare the performance of the proposed parametric bootstrap with their competitors where available, using Monte Carlo simulation studies. To maintain comparability with existing studies, we adopt part of the simulation framework of Chatterjee et al. (2008). We consider an area-level model with $\bm{x_{i}^{\prime}\beta}=0$, and five groups of small areas. Within each group, the sampling variances $D_{i}$’s remain the same. Two patterns for the $D_{i}$’s are considered: (i) $(4.0,0.6,0.5,0.4,0.2)$; (ii) $(8.0,1.2,1.0,0.8,0.4)$. For pattern (i), we take $A=1$; for pattern (ii), we take $A=2$ doubling the variances of pattern (i) while preserving the $B_{i}=D_{i}/(A+D_{i})$ ratios. To examine the effect of $m$, we consider $m=15$ and 50. With the increase of $m$, all methods improve and get closer to one another, supporting our asymptotic theory. Since we obtained virtually identical results under these two patterns, for the full study we confined attention to the pattern (i) and the results for pattern (ii) are provided in the Appendix G.
@@
$f(A)=\sum_{i=1}^{m}\frac{(y_{i}-\bm{x_{i}^{\prime}\tilde{\beta}})^{2}}{A+D_{i}}=0.$
@@ 5.1 Simulations on symmetric cases
First, we consider the scenario that $u_{i}$ is symmetrically distributed. Specifically, we assume $\mathcal{G}$ is a $t$ distribution with 9 degrees of freedom. In this setting, we compare coverage probabilities and average lengths of the following seven different prediction intervals of $\theta_{i}$:
@@
Each reported result is based on 1000 simulation runs. For all cases, we consider single bootstrap sample of size 400 and three different nominal coverages: $\alpha\in(80\%,90\%,95\%)$.
@@
Table 1: Percentages of negative estimates of $A$ ($\hat{A}^{*}$) for different estimation methods and different patterns of $D_{i}$ under $t_{9}$.
@@
SB.FH HM.FH SB.PR HM.PR FH PR DIRECT (i) $\alpha=80\%$ G1 79.94 ( 2.31 ) 80.30 ( 2.53 ) 83.7 ( 2.59 ) 84.45 ( 2.84 ) 79.89 ( 2.30 ) 78.66 ( 2.31 ) 79.43 ( 5.13 ) G2 80.15 ( 1.59 ) 80.31 ( 2.49 ) 83.63 ( 1.79 ) 84.54 ( 2.81 ) 80.38 ( 1.59 ) 81.55 ( 1.67 ) 79.85 ( 1.99 ) G3 79.71 ( 1.50 ) 79.53 ( 2.48 ) 83.12 ( 1.69 ) 83.75 ( 2.80 ) 79.86 ( 1.50 ) 80.88 ( 1.58 ) 79.99 ( 1.81 ) G4 80.47 ( 1.39 ) 80.03 ( 2.48 ) 84.06 ( 1.57 ) 83.95 ( 2.80 ) 80.49 ( 1.39 ) 82.06 ( 1.47 ) 80.29 ( 1.62 ) G5 79.65 ( 1.05 ) 80.48 ( 2.46 ) 83.32 ( 1.20 ) 84.89 ( 2.78 ) 79.99 ( 1.06 ) 81.8 ( 1.14 ) 80.32 ( 1.15 ) (i) $\alpha=90\%$ G1 89.88 ( 3.05 ) 90.35 ( 3.39 ) 94.13 ( 3.79 ) 94.25 ( 4.21 ) 88.9 ( 2.96 ) 87.82 ( 2.97 ) 89.42 ( 6.58 ) G2 90.17 ( 2.05 ) 90.37 ( 3.30 ) 93.25 ( 2.59 ) 94.27 ( 4.14 ) 89.89 ( 2.04 ) 90.58 ( 2.14 ) 90.19 ( 2.55 ) G3 90.06 ( 1.94 ) 89.34 ( 3.30 ) 93.38 ( 2.44 ) 93.76 ( 4.14 ) 90.15 ( 1.92 ) 90.78 ( 2.03 ) 90.11 ( 2.33 ) G4 90.07 ( 1.79 ) 89.50 ( 3.29 ) 93.41 ( 2.26 ) 94.24 ( 4.13 ) 90.12 ( 1.78 ) 91.19 ( 1.89 ) 90.24 ( 2.08 ) G5 90.29 ( 1.36 ) 90.28 ( 3.26 ) 93.33 ( 1.72 ) 94.2 ( 4.09 ) 90.26 ( 1.36 ) 91.56 ( 1.47 ) 90.32 ( 1.47 ) (i) $\alpha=95\%$ G1 95.13 ( 3.75 ) 95.04 ( 4.22 ) 97.3 ( 5.31 ) 97.42 ( 5.91 ) 93.57 ( 3.52 ) 92.65 ( 3.53 ) 94.77 ( 7.84 ) G2 95.11 ( 2.47 ) 95.35 ( 4.07 ) 96.99 ( 3.61 ) 97.5 ( 5.79 ) 94.87 ( 2.43 ) 95.52 ( 2.55 ) 95.38 ( 3.04 ) G3 94.96 ( 2.32 ) 94.45 ( 4.06 ) 97.05 ( 3.39 ) 97.42 ( 5.77 ) 94.70 ( 2.29 ) 95.51 ( 2.41 ) 95.00 ( 2.77 ) G4 95.06 ( 2.14 ) 95.03 ( 4.04 ) 96.86 ( 3.14 ) 97.58 ( 5.75 ) 95.12 ( 2.12 ) 95.86 ( 2.25 ) 95.07 ( 2.48 ) G5 95.31 ( 1.62 ) 94.98 ( 3.99 ) 96.87 ( 2.38 ) 97.42 ( 5.69 ) 95.28 ( 1.62 ) 96.01 ( 1.75 ) 95.17 ( 1.75 )

### v3

@@ 3 Single parametric bootstrap
$\text{E}\left[(\theta_{i}-\tilde{\theta}_{i})/\sqrt{g_{1i}}\right]^{4}$ $=$ $\frac{1}{g_{1i}^{2}}\text{E}\left[\bm{x}_{i}^{\prime}\bm{\beta}+u_{i}-(1-B_{i})y_{i}-B_{i}\bm{x}_{i}^{\prime}\bm{\beta}\right]^{4}$
$=$ $3+\frac{D_{i}^{2}}{(A+D_{i})^{2}}\left[\frac{\text{E}(u_{i}^{4})}{A^{2}}-3\right],$
@@ 4 Double parametric bootstrap
[…]
Proposition 3
@@
$\mathrm{P}(\bm{x}_{i}^{\prime}\hat{\bm{\beta}}+b_{\hat{\alpha}_{l}^{s}}\sqrt{\hat{A}}\leq\theta_{i}\leq\bm{x}_{i}^{\prime}\hat{\bm{\beta}}+b_{\hat{\alpha}_{u}^{s}}\sqrt{\hat{A}})=1-\alpha+O(m^{-1}),$ (10)
@@
See proof in the online Appendix D, which is similar to Proposition 1.
@@
While the order $O(m^{-1})$ term $T_{1}$ in (9) is theoretically positive under certain conditions, it remains unclear whether this positiveness holds when $u_{i}$ is asymmetrically distributed. Unlike Hall and Maiti (2006), we introduce a new double bootstrap method, which does not require a pivot or symmetric $u_{i}$. In general, this method reduces the coverage error to $o(m^{-1})$, even when $u_{i}$ is asymmetrically distributed. Our double bootstrap approach is based on the algorithm from Shi (1992), where the double bootstrap is proposed to obtain accurate and efficient confidence interval for parameters of interest in both nonparametric and univariate parametric distribution settings. Later on, McCullough and Vinod (1998) discussed the theory of the double bootstrap both with and without pivoting, and provided the implementations in some nonlinear production functions. In this paper, we develop the double bootstrap in the context of our mixed effect model and apply it to obtain the EBL prediction intervals of $\theta_{i}$. The framework of our double parametric bootstrap is as below:
@@
[…]
@@
      Compute $\hat{\bm{\beta}}^{*}$, $\hat{A}^{*}$, $\hat{\bm{\phi}}^{*}$ from the data $(\bm{X},\bm{Y}^{*})$, and obtain $\hat{\theta}_{i}^{*}$ and $\hat{g}_{1i}^{*}$;
@@
      Compute $\hat{\bm{\beta}}^{**}$, $\hat{A}^{**}$ from the data $(\bm{X},\bm{Y}^{**})$, also obtain $\hat{\theta}_{i}^{**}$ and $\hat{g}_{1i}^{**}$;
@@
      $\hat{G}^{*}(z)\equiv\mathrm{P}^{**}(H_{i}(\hat{\bm{\beta}}^{**},\hat{A}^{**})\leq z\,|\,\hat{\bm{\beta}}^{*},\hat{A}^{*},\hat{\phi}^{*})$ (11)
@@
      $\left\{\begin{array}[]{@{}l@{}}\mathrm{P}^{*}(H_{i}(\hat{\bm{\beta}}^{*},\hat{A}^{*})\leq q_{\hat{\alpha}_{u}}\,|\,\hat{\bm{\beta}},\hat{A},\hat{\bm{\phi}})=1-\alpha/2\\ \mathrm{P}^{**}(H_{i}(\hat{\bm{\beta}}^{**},\hat{A}^{**})\leq q_{\hat{\alpha}_{u}}\,|\,\hat{\bm{\beta}}^{*},\hat{A}^{*},\hat{\bm{\phi}}^{*})={\hat{\alpha}_{u}}.\end{array}\right$.\,, (12)
      Using the definition (11), we have $q_{\hat{\alpha}_{u}}=\hat{G}^{*-1}(\hat{\alpha}_{u})$. Then, the above system of equations is equivalent to
      $\mathrm{P}^{*}\left(H_{i}(\hat{\bm{\beta}}^{*},\hat{A}^{*})\leq\hat{G}^{*-1}(\hat{\alpha}_{u})\,|\,\hat{\bm{\beta}},\hat{A},\hat{\bm{\phi}}\right)=1-\alpha/2.$ (13)
      Rewriting (13) gives
      $\mathrm{P}^{*}\left[\mathrm{P}^{**}\left(H_{i}(\hat{\bm{\beta}}^{**},\hat{A}^{**})\leq H_{i}(\hat{\bm{\beta}}^{*},\hat{A}^{*})\,|\,\hat{\bm{\beta}}^{*},\hat{A}^{*},\hat{\bm{\phi}}^{*}\right)\leq\hat{\alpha}_{u}\,|\,\hat{\bm{\beta}},\hat{A},\hat{\bm{\phi}}\right]=1-\alpha/2.$ (14)
@@
      We will use $Z_{j}$ to adjust the first-stage intervals. After all bootstrapping operations are complete, we have estimates $Z_{j}$, $j=1,2,\cdots,J$, where $J$ is the number of first-stage bootstrap operations. Sort $Z_{j}$ and choose the $1-\alpha/2$ quantile as the percentile point $\hat{\alpha}_{u}$ for defining the double-bootstrap upper limit.
@@
      $\mathrm{P}^{*}(H_{i}(\hat{\bm{\beta}}^{*},\hat{A}^{*})\leq q_{\alpha_{u}}\,|\,\hat{\bm{\beta}},\hat{A},\hat{\bm{\phi}})=\hat{\alpha}_{u}.$ (15)
@@
The algorithm above shows that the single parametric bootstrap $\hat{I}^{s}_{i}$ is calibrated by the second-stage bootstrap. Such calibration improves the coverage accuracy to $o(m^{-1})$ even when the pivot does not exist. One more advantage of our double bootstrap algorithm is that it avoids the problem of over correction and so make it more practical than that of Hall and Maiti (2006).
Theorem 2
@@
$\mathrm{P}\left(\hat{\theta}_{i}+q_{\hat{\alpha}^{d}_{l}}\sqrt{\hat{g}_{1i}}\leq\theta_{i}\leq\hat{\theta}_{i}+q_{\hat{\alpha}_{u}^{d}}\sqrt{\hat{g}_{1i}}\right)=1-\alpha+o(m^{-1}),$ (16)
@@
See proof of the theorem in online Appendix E.
@@ 5 Monte Carlo Simulations
In this section, we compare the performance of the proposed parametric bootstrap with their competitors where available, using Monte Carlo simulation studies. To maintain comparability with existing studies, we adopt part of the simulation framework of Chatterjee et al. (2008). We consider an area-level model with $\bm{x}_{i}^{\prime}\bm{\beta}=0$, and five groups of small areas. Within each group, the sampling variances $D_{i}$’s remain the same. Two patterns for the $D_{i}$’s are considered: (i) $(4.0,0.6,0.5,0.4,0.2)$; (ii) $(8.0,1.2,1.0,0.8,0.4)$. For pattern (i), we take $A=1$; for pattern (ii), we take $A=2$ doubling the variances of pattern (i) while preserving the $B_{i}=D_{i}/(A+D_{i})$ ratios. To examine the effect of $m$, we consider $m=15$ and 50. With the increase of $m$, all methods improve and get closer to one another, supporting our asymptotic theory. Since we obtained virtually identical results under these two patterns, for the full study we confined attention to the pattern (i) and the results for pattern (ii) are provided in the Supplementary Material.
@@
$f(A)=\sum_{i=1}^{m}\frac{(y_{i}-\bm{x}_{i}^{\prime}\tilde{\bm{\beta}})^{2}}{A+D_{i}}=0.$
@@ 5.1 Simulations on symmetric cases
First, we consider the scenario that the random effects $u_{i}$ are symmetrically distributed. Specifically, we assume $\mathcal{G}$ is a $t$ distribution with 9 degrees of freedom. In this setting, we compare coverage probabilities and average lengths of the following seven different prediction intervals of $\theta_{i}$:
@@
Each reported result is based on 1000 simulation runs. For all cases, we consider single bootstrap sample of size 400 and three different nominal coverages: $\alpha\in(80\%,90\%,95\%)$. All computations are carried out in R on a 13th Gen Intel(R) Core(TM) i7-13700F 2.10 GHz processor. The average computation time per simulation run for the proposed single bootstrap method is 10.05 seconds for $m=15$ and 11.55 seconds for $m=50$.
@@
Table 1: Percentages of negative estimates of AA ($\hat{A}^{*}$) for different estimation methods and different patterns of $D_{i}$ under $t_{9}$.
@@
SB.FH HM.FH SB.PR HM.PR FH PR DIRECT (i) $\alpha=80\%$ G1 79.94 ( 2.31 ) 80.30 ( 2.53 ) 83.70 ( 2.59 ) 84.45 ( 2.84 ) 79.89 ( 2.30 ) 78.66 ( 2.31 ) 79.43 ( 5.13 ) G2 80.15 ( 1.59 ) 80.31 ( 2.49 ) 83.63 ( 1.79 ) 84.54 ( 2.81 ) 80.38 ( 1.59 ) 81.55 ( 1.67 ) 79.85 ( 1.99 ) G3 79.71 ( 1.50 ) 79.53 ( 2.48 ) 83.12 ( 1.69 ) 83.75 ( 2.80 ) 79.86 ( 1.50 ) 80.88 ( 1.58 ) 79.99 ( 1.81 ) G4 80.47 ( 1.39 ) 80.03 ( 2.48 ) 84.06 ( 1.57 ) 83.95 ( 2.80 ) 80.49 ( 1.39 ) 82.06 ( 1.47 ) 80.29 ( 1.62 ) G5 79.65 ( 1.05 ) 80.48 ( 2.46 ) 83.32 ( 1.20 ) 84.89 ( 2.78 ) 79.99 ( 1.06 ) 81.8 ( 1.14 ) 80.32 ( 1.15 ) (i) $\alpha=90\%$ G1 89.88 ( 3.05 ) 90.35 ( 3.39 ) 94.13 ( 3.79 ) 94.25 ( 4.21 ) 88.90 ( 2.96 ) 87.82 ( 2.97 ) 89.42 ( 6.58 ) G2 90.17 ( 2.05 ) 90.37 ( 3.30 ) 93.25 ( 2.59 ) 94.27 ( 4.14 ) 89.89 ( 2.04 ) 90.58 ( 2.14 ) 90.19 ( 2.55 ) G3 90.06 ( 1.94 ) 89.34 ( 3.30 ) 93.38 ( 2.44 ) 93.76 ( 4.14 ) 90.15 ( 1.92 ) 90.78 ( 2.03 ) 90.11 ( 2.33 ) G4 90.07 ( 1.79 ) 89.50 ( 3.29 ) 93.41 ( 2.26 ) 94.24 ( 4.13 ) 90.12 ( 1.78 ) 91.19 ( 1.89 ) 90.24 ( 2.08 ) G5 90.29 ( 1.36 ) 90.28 ( 3.26 ) 93.33 ( 1.72 ) 94.2 ( 4.09 ) 90.26 ( 1.36 ) 91.56 ( 1.47 ) 90.32 ( 1.47 ) (i) $\alpha=95\%$ G1 95.13 ( 3.75 ) 95.04 ( 4.22 ) 97.30 ( 5.31 ) 97.42 ( 5.91 ) 93.57 ( 3.52 ) 92.65 ( 3.53 ) 94.77 ( 7.84 ) G2 95.11 ( 2.47 ) 95.35 ( 4.07 ) 96.99 ( 3.61 ) 97.5 ( 5.79 ) 94.87 ( 2.43 ) 95.52 ( 2.55 ) 95.38 ( 3.04 ) G3 94.96 ( 2.32 ) 94.45 ( 4.06 ) 97.05 ( 3.39 ) 97.42 ( 5.77 ) 94.70 ( 2.29 ) 95.51 ( 2.41 ) 95.00 ( 2.77 ) G4 95.06 ( 2.14 ) 95.03 ( 4.04 ) 96.86 ( 3.14 ) 97.58 ( 5.75 ) 95.12 ( 2.12 ) 95.86 ( 2.25 ) 95.07 ( 2.48 ) G5 95.31 ( 1.62 ) 94.98 ( 3.99 ) 96.87 ( 2.38 ) 97.42 ( 5.69 ) 95.28 ( 1.62 ) 96.01 ( 1.75 ) 95.17 ( 1.75 )